OpenStudy (mathhelp346):

How can I turn word problems into algebra equations? Can you put in kid words too? :}

OpenStudy (mathhelp346):

Anyoneee?

OpenStudy (ybarrap):

Assign to each unknown a variable to represent it Take the sentence and reduce the relationships to less complex language continue and continue to reduce the number of words until you're left with just the variables you've defined you should end up with just variables, some equal signs and operators (+,-,/,*)

OpenStudy (mathhelp346):

???

OpenStudy (anonymous):

do you have a specific example?

OpenStudy (mathhelp346):

...no

OpenStudy (anonymous):

would you like me to make one up?

OpenStudy (mathhelp346):

sure

OpenStudy (anonymous):

ok together our ages total to 62 i am 12 years older than you are. how old am i?

OpenStudy (mathhelp346):

50

OpenStudy (mathhelp346):

it's wrong

OpenStudy (anonymous):

so this is what we do

OpenStudy (anonymous):

first introduce a variable, you can pick any letter. then write down precisely what it means. make sure you know it is going to be a number. i will use "m"

OpenStudy (anonymous):

and say that "m is my age"

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

then i will see if i can find some other expression involving only m. for example, if i am 40 then you must be 30 (because i am ten years older) and if i am 35 you must be 25 for the same reason

OpenStudy (anonymous):

so if my age is "m" then your age must be \[m-10\]

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

so now i have both our ages written as variables. mine is \[m\] and yours is \[m-10\]

OpenStudy (mathhelp346):

ok

OpenStudy (ybarrap):

sage = 12 + you, sat said his age is 12 more than yours you + sage = 62, he said that your age and his combined to 62 you = ?, what is your age? you = 62 - sage, your age from the 2nd equation above is this = 62 - (12 + you), substitute sage's age in equation 1 here to have only on variable, "you" you + you = 62 - 12, now solve by arranging terms 2you = 50, solve some more you = 50/2, arithmetic, easy = 25 That's it, turn language into variables into equations.

OpenStudy (anonymous):

finally i find another expression for our ages. i know that the total of our ages is 62, so i write that down as well

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

my age plus your age is 62 \[m+m-10=62\]

OpenStudy (anonymous):

now i have some algebra that i can solve with no words attached

OpenStudy (mathhelp346):

wait where did the second m come from

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

my age is \[m\] and your age is \[m-10\] right?

OpenStudy (mathhelp346):

yea i understand now

OpenStudy (anonymous):

so when i add them together i get \[m+m-10=62\] which says my age plus your age is 62

OpenStudy (mathhelp346):

well i mean i understand where you got the m from

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

if parentheses help i could put \[(m)+(m-10)=62\] but they are unnecessary

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

now it is customary to write \[2m\] instead of \[m+m\]so we could rewrite this as \[2m-10=62\]

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

to solve now takes two steps. add ten to both sides \[2m-10+10=62+10\] \[2m=72\]

OpenStudy (mathhelp346):

why do you need to add 10?

OpenStudy (anonymous):

then to find m divide both sides by two or just think "twice my numbers is 72, so my number must be half of 72 and half of 72 is 36

OpenStudy (anonymous):

why did i add 10?

OpenStudy (anonymous):

i have an equation that is \[2m-10=62\] and i am trying to find the number m

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

in other words my job is to get \[m=\text{something}\]

OpenStudy (mathhelp346):

so you're trying to get rid of the 10?

OpenStudy (anonymous):

exactly. i want m all alone, so i have to get rid of the ten first

OpenStudy (anonymous):

think of it this way. i have \[2m-10=62\] and i say "what was this before i subtracted 10?" well it must have been 72 first right? i know this because \[62+10=72\]so i write \[2m-10=62\] \[2m=72\]

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

then i know that two times my number is 72, so my number must be half of 72 and half of 72 is 36

OpenStudy (mathhelp346):

but if you subtracted 10, wouldn't it be \[2m = 62\]

OpenStudy (anonymous):

hold on we started with \[2m-10=62\] right?

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

so is it clear that if \[2m-10=62\] then \[2m=72\]?

OpenStudy (mathhelp346):

...not really

OpenStudy (anonymous):

in other words i have a quantity 2m when i subtract ten from it i get 62 that means the quantity itself must be 72

OpenStudy (anonymous):

lets think like this. i have some money in my pocket. i spend 10 dollars and am left with $62 who much money did i start with?

OpenStudy (mathhelp346):

ohhh ok u started with $72

OpenStudy (anonymous):

right. like saying \[\text{my wallet} - 10 = 62\]

OpenStudy (anonymous):

so i had $72 in my wallet to begin with

OpenStudy (anonymous):

and you know this because you added 10 to 62 to get 72

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

in our algebra problem we have \[2m-10=62\] so we know that \[2m=72\]

OpenStudy (anonymous):

and if twice my age is 72 then my age must be 36

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

in math we would say "divide both sides by 2" but it is really just common sense. two times my age is 72, my age is half of 72 which is 36

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

so now you know my age. my age is 36 and we also know your age. you are ten years younger than i am so you must be 26

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

and you see that we have solved it because the total of our ages is \[36+26=62\] and the numbers are ten apart

OpenStudy (mathhelp346):

yep

OpenStudy (anonymous):

if you have a specific problem like this feel free to post it and you will get lots of answers

OpenStudy (mathhelp346):

thank you soooooooooooooo much satellite